# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 5.3. The A-B-C-D Lemma and Criteria for the regular points
Taking the divergence of (3.3)${}_{1}$ yields
(5.26) $-\Delta{\pi}$ $=\mathop{\rm div}\nolimits^{2}\Big{[}{{\bf{v}}}\otimes{{\bf{v}}}+\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}(\nabla{{\bf{d}}}:\nabla{{\bf{d}}}){\rm I}_{3}+{{\bf{z}}}\otimes{{\bf{v}}}+{{\bf{v}}}\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big{]}$
$=\mathop{\rm div}\nolimits^{2}\Big{[}({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}})_{\rho}+(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}})_{\rho}{\rm I}_{3}$
$-\frac{1}{2}(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$+{{\bf{z}}}\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big{]}$
$:=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho}.$
@@
$-\Delta\widecheck{\pi}=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho/2},$
@@
$-\Delta({\pi}-\widecheck{\pi})=0\text{ in }B_{\rho/2}.$
@@
$\displaystyle\int_{\mathbb{T}^{3}}|\widecheck{\pi}|^{\frac{3}{2}}{\,\mathrm{d}x}$ $\displaystyle\lesssim\int_{\mathbb{T}^{3}}\eta^{3}\left(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}\right)$
@@
$\displaystyle\lesssim\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{B_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+1)$
Notice that ${\pi}-\widecheck{\pi}$ is harmonic in $B_{\rho/2}$, and we have that for $2\displaystyle 0<r\leq\rho/$ 2,
@@
$\displaystyle\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}}$ $\displaystyle\lesssim\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{\pi}|^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{2}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim\left(\frac{r}{\rho}\right)D(\rho)+\left(\frac{\rho}{r}\right)^{2}\Big{[}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}$
$\displaystyle+\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}}(t))_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}+\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\rho^{3}\Big{]}.$
@@
$\displaystyle\lesssim\left(\sup_{-\rho^{2}\leq t\leq 0}\frac{1}{\rho}\int_{B_{\rho}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{\mathcal{Q}_{\rho}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}=A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and by Hölder’s inequality and Poincaré’s inequality, we can show
@@
$\displaystyle\lesssim\rho^{2}\left(\sup_{-\rho^{2}\leq t\leq 0}\int_{B_{\rho}}|\nabla{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{Q_{\rho}}|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\leq A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
@@
Now we have the interpolation estimate (5.24) and the pressure estimate (5.25), the proof of Theorem 5.5 is based on a dichotomy argument:
Proof of Theorem 5.5.
@@
Applying the local energy inequality in Lemma 4.1 with $\varphi^{2}$, the $L^{\infty}$-bounded of $({{\bf{z}}},{{\bf{d}}})$, and integrating by parts, we obtain that
@@
$\displaystyle\lesssim\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})(|\partial_{t}\varphi|+|\nabla\varphi|^{2}+|\nabla^{2}\varphi|)$
@@
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+|\nabla{\bf{d}}|^{2}|{{\bf{v}}}||\nabla\varphi|+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}|{{\bf{z}}}|^{2}|\nabla\varphi|$
@@
$\displaystyle=\sup_{-\left(\frac{\theta\rho}{2}\right)^{2}\leq t\leq 0}\frac{2}{\theta\rho}\int_{B_{\frac{\theta\rho}{2}}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+\frac{2}{\theta\rho}\int_{\mathcal{Q}_{\frac{\theta\rho}{2}}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\displaystyle\lesssim\sup_{-(\theta\rho)^{2}\leq t\leq 0}\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
@@
$\displaystyle A(\frac{1}{2}\theta\rho)+B(\frac{1}{2}\theta\rho)$ $\displaystyle\lesssim\left[C^{\frac{2}{3}}(\theta\rho)+A^{\frac{1}{2}}(\theta\rho)B^{\frac{1}{2}}(\theta\rho)C^{\frac{1}{3}}(\theta\rho)+C^{\frac{1}{3}}(\theta\rho)D^{\frac{2}{3}}(\theta\rho)+(\theta\rho)^{3}\right]$
$\displaystyle\lesssim C^{\frac{2}{3}}(\theta\rho)+A(\theta\rho)B(\theta\rho)+D^{\frac{4}{3}}(\theta\rho)+(\theta\rho)^{3}.$
@@
On the other hand, from Lemma 5.7 and 5.8 we can conclude that
@@
For $\varepsilon_{0}$ given by Lemma 5.2, let $\theta_{0}\in(0,\frac{1}{2})$ be such that
@@
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Furthermore, we can obtain from (5.24) that
@@
holds for all $\rho\in(0,\rho_{0})$ provided $\varepsilon_{1}=\varepsilon_{1}(\theta_{0},\varepsilon_{0})>0$ is chosen sufficiently small. Hence there exists a $k_{0}>0$ such that
@@
which satisfies the assumption of Lemma 5.2, and as a consequence, $z_{0}=0$ is a regular point. ∎
5.4. The Hausdorff measure estimate
Now we have all the ingredients to prove Theorem 3.3.
Proof of Theorem 3.3.
If $z\in\Sigma$, then by Theorem 5.5,
@@ 5.4. The Hausdorff measure estimate
$\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
@@
(5.27) $\mathcal{P}_{5\eta}^{1}(\Sigma)$ $\leq\sum_{i=1}^{\infty}5r_{i}\leq\frac{5}{\varepsilon_{1}}\sum_{i=1}^{\infty}\int_{\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$

### v2

@@ 5.3 The A-B-C-D Lemma and Criteria for the regular points
Taking the divergence of (11)1 yields
$\displaystyle-\Delta{\pi}=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\Big[{{\bf{v}}}\otimes{{\bf{v}}}+\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}(\nabla{{\bf{d}}}:\nabla{{\bf{d}}}){\rm I}_{3}+{{\bf{z}}}\otimes{{\bf{v}}}+{{\bf{v}}}\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big]$
$\displaystyle=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\Big[({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$\displaystyle+(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})\odot(\nabla{{\bf{d}}})_{\rho}+(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho})$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}):(\nabla{{\bf{d}}})_{\rho}{\rm I}_{3}$
$\displaystyle-\frac{1}{2}(\nabla{{\bf{d}}})_{\rho}\odot(\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}){\rm I}_{3}$
$\displaystyle+{{\bf{z}}}\otimes({{\bf{v}}}-({{\bf{v}}})_{\rho})+({{\bf{v}}}-({{\bf{v}}})_{\rho})\otimes{{\bf{z}}}+{{\bf{z}}}\otimes{{\bf{z}}}-F({{\bf{d}}}){\rm I}_{3}\Big]$
$\displaystyle:=$ $\displaystyle\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho}.$
@@
$\Delta\widecheck{\pi}=\mathop{\rm div}\nolimits^{2}\mathcal{G}\text{ in }B_{\rho/2},$
@@
$\Delta({\pi}-\widecheck{\pi})=0\text{ in }B_{\rho/2}.$
@@
$\displaystyle\int_{\mathbb{T}^{3}}|\widecheck{\pi}|^{\frac{3}{2}}{\,\mathrm{d}x}\lesssim$ $\displaystyle\int_{\mathbb{T}^{3}}\eta^{3}\left(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}\right)$
@@
$\displaystyle\lesssim$ $\displaystyle\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3})+|(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}\int_{B_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle+\int_{B_{\rho}}(|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+1)$
Notice that ${\pi}-\widecheck{\pi}$ is harmonic in $B_{\rho/2}$, and we have that for $0<r\leq\rho/2$,
@@
$\displaystyle\frac{1}{r^{2}}\int_{\mathcal{Q}_{r}(0)}|{\pi}|^{\frac{3}{2}}\lesssim$ $\displaystyle\left(\frac{r}{\rho}\right)\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{\pi}|^{\frac{3}{2}}+\left(\frac{\rho}{r}\right)^{2}\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|\widecheck{\pi}|^{\frac{3}{2}}$
$\displaystyle\lesssim$ $\displaystyle\left(\frac{r}{\rho}\right)D(\rho)+\left(\frac{\rho}{r}\right)^{2}\Big[\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{3}+|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{3}$
$\displaystyle+\frac{1}{\rho^{2}}(\sup_{-\rho^{2}\leq t\leq 0}|(\nabla{{\bf{d}}}(t))_{\rho}|^{\frac{3}{2}})\int_{\mathcal{Q}_{\rho}}|\nabla{{\bf{d}}}-(\nabla{{\bf{d}}})_{\rho}|^{\frac{3}{2}}$
$\displaystyle+\frac{1}{\rho^{2}}\int_{\mathcal{Q}_{\rho}}|{{\bf{v}}}-({{\bf{v}}})_{\rho}|^{\frac{3}{2}}+\rho^{3}\Big].$
@@
$\displaystyle\lesssim\left(\sup_{-\rho^{2}\leq t\leq 0}\frac{1}{\rho}\int_{B_{\rho}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{\mathcal{Q}_{\rho}}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\right)^{\frac{3}{4}}$
$\displaystyle=A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and by Hölder’s inequality and Poincaré’s inequality, we can verify that
@@
$\displaystyle\lesssim$ $\displaystyle\rho^{2}\left(\sup_{-\rho^{2}\leq t\leq 0}\int_{B_{\rho}}|\nabla{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}\left(\frac{1}{\rho}\int_{Q_{\rho}}|\nabla^{2}{{\bf{d}}}|^{2}\right)^{\frac{3}{4}}$
$\displaystyle\leq$ $\displaystyle A^{\frac{3}{4}}(\rho)B^{\frac{3}{4}}(\rho),$
and
@@
Now we have the interpolation estimate (74) and the pressure estimate (75), the proof of Theorem 8 is based on a dichotomy argument:
Proof of Theorem 8.
@@
Applying the local energy inequality in (15) with $\varphi^{2}$, the $L^{\infty}$-bounded of $({{\bf{z}}},{{\bf{d}}})$, and integrating by parts, we obtain that
@@
$\displaystyle\lesssim$ $\displaystyle\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})(|\partial_{t}\varphi|+|\nabla\varphi|^{2}+|\nabla^{2}\varphi|)$
@@
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+|\nabla{\bf{d}}|^{2}|{{\bf{v}}}||\nabla\varphi|$
$\displaystyle+\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}|{{\bf{z}}}|^{2}|\nabla\varphi|$
@@
$\displaystyle=$ $\displaystyle\sup_{-\left(\frac{\theta\rho}{2}\right)^{2}\leq t\leq 0}\frac{2}{\theta\rho}\int_{B_{\frac{\theta\rho}{2}}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})+\frac{2}{\theta\rho}\int_{\mathcal{Q}_{\frac{\theta\rho}{2}}(0)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$
$\displaystyle\lesssim$ $\displaystyle\sup_{-(\theta\rho)^{2}\leq t\leq 0}\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}}(|{{\bf{v}}}|^{2}+|\nabla{{\bf{d}}}|^{2})\varphi^{2}+\frac{1}{\theta\rho}\int_{\mathbb{T}^{3}\times[-(\theta\rho)^{2},0]}(|\nabla{{\bf{u}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\varphi^{2}$
@@
$\displaystyle A(\frac{1}{2}\theta\rho)+B(\frac{1}{2}\theta\rho)$
$\displaystyle\lesssim$ $\displaystyle\left[C^{\frac{2}{3}}(\theta\rho)+A^{\frac{1}{2}}(\theta\rho)B^{\frac{1}{2}}(\theta\rho)C^{\frac{1}{3}}(\theta\rho)+C^{\frac{1}{3}}(\theta\rho)D^{\frac{2}{3}}(\theta\rho)+(\theta\rho)^{3}\right]$
$\displaystyle\lesssim$ $\displaystyle C^{\frac{2}{3}}(\theta\rho)+A(\theta\rho)B(\theta\rho)+D^{\frac{4}{3}}(\theta\rho)+(\theta\rho)^{3}.$
@@
On the other hand, from Lemma 10 and 11 we can conclude that
@@
For $\varepsilon_{0}$ given by Lemma 5, let $\theta_{0}\in(0,\frac{1}{2})$ be such that
@@
holds for all $0<\rho<\rho_{0}$ and $k\geq 1$. Furthermore, we can obtain from (74) that
@@
holds for all $\rho\in(0,\rho_{0})$ provided that $\varepsilon_{1}=\varepsilon_{1}(\theta_{0},\varepsilon_{0})>0$ is chosen sufficiently small. Hence there exists a $k_{0}>0$ such that
@@
which satisfies the assumption of Lemma 5, and as a consequence, $z_{0}=0$ is a regular point. ∎
5.4 The Hausdorff measure estimate
Now we have all the ingredients to prove Theorem 2. It is based on a standard covering argument which we detailed below for the reader’s convenience.
Proof of Theorem 2.
If $z\in\Sigma$, then by Theorem 8,
@@ 5.4 The Hausdorff measure estimate
$\frac{1}{r}\int_{\mathcal{Q}_{r}(z)}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})\geq\varepsilon_{1}.$
@@
$\mathcal{P}_{5\eta}^{1}(\Sigma)$ $\leq\sum_{i=1}^{\infty}5r_{i}\leq\frac{5}{\varepsilon_{1}}\sum_{i=1}^{\infty}\int_{\mathcal{Q}_{r_{i}}(z_{i})}(|\nabla{{\bf{v}}}|^{2}+|\nabla^{2}{{\bf{d}}}|^{2})$ (76)